# Gelation in Vector Multiplicative Coalescence and Extinction in Multi-Type Poisson Branching Processes

### v1

@@ Abstract.
Random coalescent processes and branching processes are two fundamental constructs in the field of stochastic processes, each with a rich history and a wide range of applications. Though developed within distinct contexts, in this note we present a novel connection between a multi-type (vector) multiplicative coalescent process and a multi-type branching process with Poisson offspring distributions. More specifically, we show that the equations that govern the phenomenon of gelation in the vector multiplicative coalescent process are equivalent to the set of equations that yield the extinction probabilities of the corresponding multi-type Poisson branching process. We then leverage this connection with two applications, one in each direction. The first is a new quick proof of gelation in the vector multiplicative coalescent process using a well known result of branching processes, and the second is a new series expression for the extinction probabilities of the multi-type Poisson branching process using results derived from the theory of vector multiplicative coalescence. While the correspondence is fairly straightforward, it illuminates a deep connection between these two paradigms which we hope will continue to reveal new insights and potential for cross-disciplinary research.
@@
Vector-multiplicative coalescent, Gelation, Multitype branching processes, Extinction, Lambert-Euler inversion
1. Vector Multiplicative Coalescent Processes
In this section, we give an overview of the theory of vector multiplicative coalescent processes. The complete theory was derived by two of the authors in [13]. Following the convention in [13], to define the vector multiplicative coalescent process, we use bra-ket notation. Specifically, $|{\bf x}\rangle$ will denote the column vector representation of vector ${\bf x}\in\mathbb{R}^{k}$, and $\langle{\bf x}|$ will denote the row vector representation of vector ${\bf x}\in\mathbb{R}^{k}$. For $c\in\mathbb{R}$ and ${\bf x}\in\mathbb{R}^{k}$, $\,c|{\bf x}\rangle$ will represent the product $c{\bf x}$, a column vector. Correspondingly, $\langle{\bf x}|{\bf y}\rangle=\langle{\bf y}|{\bf x}\rangle$ will be the dot product of ${\bf x}$ and ${\bf y}$ in $\mathbb{R}^{k}$. Finally, for a matrix $M\in\mathbb{R}^{k\times k}$, $\langle{\bf x}|M|{\bf y}\rangle$ will represent the product ${\bf x}^{\sf T}M{\bf y}$ resulting in a scalar.
@@ 1. Vector Multiplicative Coalescent Processes
with initial conditions $\zeta_{\bf x}(0)=\alpha_{i}$ if $x={\bf e}_{i}$, and $\zeta_{\bf x}(0)=0$ otherwise. These equations describe the evolution of the size distribution of clusters in a system where clusters merge multiplicatively as defined in (1). More specifically, the first term on the right hand side in (2) corresponds to the depletion of clusters of size x after merging with other clusters and the second term corresponds to coagulation of smaller clusters to produce one of size x. For further details on Smoluchowski equations in various settings including applications, we refer the reader to [17, 15, 4, 12, 6].
@@
with the same initial conditions coincide as long as the second order moments are convergent. The unique solution $\zeta_{\bf x}(t)$ of the above modified Smoluchowski equations was derived in [13]. Specifically, for a vector ${\bf x}\in\mathbb{Z}_{+}^{k}$ let $\,{\bf x}!=x_{1}!x_{2}!\ldots x_{k}!\,$ and for vectors ${\bf a}$ and ${\bf b}$ in $\mathbb{R}^{k}$ let $\,{\bf a}^{\bf b}=a_{1}^{b_{1}}a_{2}^{b_{2}}\ldots a_{k}^{b_{k}}\,$ whenever $a_{i}^{b_{i}}$ is uniquely defined for all $i$. Now, consider a complete graph $K_{k}$ consisting of vertices $\{1,\ldots,k\}$ with weights $w_{i,j}=w_{j,i}\geq 0$ assigned to its edges $[i,j]$ ($i\not=j$). Let the weight $W(\mathcal{T})$ of a spanning tree $\mathcal{T}$ be the product of the weights of all of its edges. Finally, let $\tau(K_{k},w_{i,j})=\sum\limits_{\mathcal{T}}W(\mathcal{T})$ denote the weighted spanning tree enumerator, i.e., the sum of weights of all spanning trees in $K_{k}$. In terms of these notations, the solution to the modified Smoluchowski equations that yield the limiting fraction of clusters of size ${\bf x}$ is given by
@@
1.1. Random graphs and an alternative proof of equation (4)
The following alternative proof of (4) is of independent interest, as it uses the connection between the vector multiplicative coalescent process $C$ and the theory of random graphs. This connection works as follows: clusters of particles in $C$ correspond to small connected components of a certain random graph $G$. The so-called giant component of $G$ corresponds to the gel in $C$ for details, see the next section. The correspondence between $C$ and $G$ was first described in [13], to which we refer the reader for a full account. Here, we include only enough detail to describe the proof of (4).
@@ 1.1. Random graphs and an alternative proof of equation (4)
With $\alpha$ and $V$ fixed, there will be a critical value $t^{\prime}$ of $t$ below which there is almost surely no giant component, and above which there is almost surely a giant component. Writing $A={\rm diag}({\boldsymbol{\alpha}})$, this critical value $t^{\prime}$ is such that the largest eigenvalue of the matrix $t^{\prime}VA$ equals 1 (i.e., $t^{\prime}$ is the reciprocal of the spectral radius of $VA$).
@@
For simplicity, we first address the case where $k=\alpha_{1}=v_{11}=1$. In this case, setting ${\bf x}=k$, we have
(5) $\zeta_{k}(t)=\phi_{k}(t)=\frac{k^{k-2}t^{k-1}e^{-kt}}{k!}$
The random graphs proof of this is as follows. We calculate $p_{k}(t)=k\phi_{k}(t)$, which is the probability that a given vertex lies in a connected component $G_{k}$ of size $k$ (which, with probability tending to 1, is a tree - see [5] for a proof). Now, setting $p=t/n$ for the edge probability, we see that:
  • There are $\binom{n-1}{k-1}$ choices for the other vertices in $G_{k}$
  • There are $k^{k-2}$ labeled trees on these chosen vertices
  • Each tree is present with probability $p^{k-1}$
  • The edges between $G_{k}$ and the other vertices are missing with probability $(1-p)^{k(n-k)}$
@@
$p_{k}(t)$ $\sim\binom{n-1}{k-1}\cdot k^{k-2}\cdot p^{k-1}\cdot(1-p)^{k(n-k)}\sim\frac{n^{k-1}}{(k-1)!}\cdot k^{k-2}\cdot\left(\frac{t}{n}\right)^{k-1}\cdot e^{-pkn}$
$=\frac{1}{(k-1)!}\cdot k^{k-2}\cdot t^{k-1}\cdot e^{-tk}=\frac{k^{k-1}t^{k-1}e^{-kt}}{k!},$
which, recalling that $p_{k}(t)=k\phi_{k}(t)$, establishes (5).
Next we consider the general case. For this we once again require the weighted spanning tree enumerator defined above, although here it arises in a slightly different form. Write $K_{\bf x}(V)$ for the complete graph on $|{\bf x}|=\sum_{i}x_{i}$ vertices with $x_{i}$ vertices of type $i$, and where the weight of an edge between vertices of types $i$ and $j$ is $v_{ij}$. Suppressing the dependence on $V$, write $T_{\bf x}$ for the sum of the weights of all the spanning trees of $K_{\bf x}(V)$. Then $T_{\bf x}=\tau(K_{k},x_{i}x_{j}v_{i,j})$, so that the two definitions of the weighted spanning tree enumerator agree.
@@
  • These form a tree with probability $T_{\bf x}\cdot(t/n)^{|{\bf x}|-1}$
@@
$p_{\bf x}(t)$ $\sim\sum_{i}x_{i}\frac{n^{|{\bf x}|-1}\alpha_{1}^{x_{1}}\cdots\alpha_{k}^{x_{k}}}{x_{1}!\cdots x_{k}!}\cdot T_{\bf x}\cdot\left(\frac{t}{n}\right)^{|{\bf x}|-1}\cdot{\rm exp}\left(-\sum_{i}\sum_{j}x_{i}v_{ij}\alpha_{j}t\right)$
$=\sum_{i}x_{i}\frac{\alpha_{1}^{x_{1}}\cdots\alpha_{k}^{x_{k}}}{x_{1}!\cdots x_{k}!}\cdot T_{\bf x}\cdot t^{|{\bf x}|-1}\cdot{\rm exp}\left(-\sum_{i}\sum_{j}x_{i}v_{ij}\alpha_{j}t\right),$
@@
$\zeta_{\bf x}(t)=\phi_{\bf x}(t)=\frac{\alpha_{1}^{x_{1}}\cdots\alpha_{k}^{x_{k}}}{x_{1}!\cdots x_{k}!}\cdot T_{\bf x}\cdot t^{|{\bf x}|-1}\cdot{\rm exp}\left(-\sum_{i}\sum_{j}x_{i}v_{ij}\alpha_{j}t\right),$
@@
1.2. Gelation and Lamber-Euler Inversion
It is natural to ask if the total mass for a system described by the Smoluchowski coagulation model is conserved. In [11, 16, 18, 19], it was shown that for certain classes of coagulation kernels, this is not always possible as a system can undergo a phase transition. For example, Hendriks et al. [9] studied this phase transition for the model with the multiplicative kernel (in one-dimensional setting), and they showed that the total mass decreases after a finite time, observing a surprising phenomenon known as gelation. Broadly speaking, gelation is the break down of conservation of mass in the system and can be physically interpreted as being caused by the formation of a giant cluster called the gel.
For the coalescent process defined in section (1), the initial total mass vector is assumed to be $\,\sum_{\bf x}\zeta_{\bf x}(0)|{\bf x}\rangle=|\boldsymbol{\alpha}\rangle$. Therefore, we define the gelation as the loss of total mass $\,\sum_{\bf x}\zeta_{\bf x}(t)|{\bf x}\rangle\,$ of the system after a critical time $T_{gel}$ called gelation time. That is
@@ 1.2. Gelation and Lamber-Euler Inversion
where for a vector ${\bf x}\in\mathbb{R}^{k}$ with coordinates $x_{i}$, $D[x_{i}]$ denotes the diagonal matrix with entries $x_{i}$, and $\rho(A)$ denotes the spectral radius of matrix $A$.
@@
The following theorem was proved in [13].
@@
Finally, in [13], it was shown that the total mass of the vector multiplicative coalescent process could be expressed in terms of the solution of the multidimensional Lambert-Euler inversion problem as the following:
@@
Then Theorem 1.2 proves the existence of the gelation phenomenon for the vector multiplicative coalescent process and the gelation time given in (7).
@@ 2. Multi-Type Poisson Branching Processes
In other words, the entry $m_{ij}$ represents the average number of offspring of type $j$ produced by an individual of type $i$ in one generation.
@@ 2.1. Extinction Probability
As gelation is to vector multiplicative coalescent processes, a fundamental question in the study of branching processes is whether the process becomes extinct or not as the number of generations $n$ tends to infinity. Recall that in the classical case, if the expected number of the offspring distribution is less than (or equal to) 1, then eventually the population dies out. But if the expected number is greater than 1, then the probability of extinction is the smallest non-negative solution to the equation $s=\varphi(s)$, where $\varphi$ is the probability generating function of the corresponding offspring distribution.

### v2

@@ Abstract.
In this note, we present a novel connection between a multi-type (vector) multiplicative coalescent process and a multi-type branching process with Poisson offspring distributions. More specifically, we show that the equations that govern the phenomenon of gelation in the vector multiplicative coalescent process are equivalent to the set of equations that yield the extinction probabilities of the corresponding multi-type Poisson branching process. We then leverage this connection with two applications, one in each direction. The first is a new quick proof of gelation in the vector multiplicative coalescent process, and the second is a new series expression for the extinction probabilities of the multi-type Poisson branching process.
@@
Vector-multiplicative coalescent, Gelation, Multitype branching processes, Extinction, Lambert-Euler inversion, Random graphs
1. Introduction
The main motivation for this work is paper [16] where the authors used the multidimensional Lambert-Euler inversion in the analysis of the vector multiplicative coalescent process. More specifically, they solved the multidimensional Lambert-Euler inversion and utilized the solution to investigate the gelation phenomenon in the coagulation model with the vector-multiplicative kernel. Our goal in this note is to present an interesting connection between the vector multiplicative coalescent process and a multi-type branching process with Poisson offspring distributions. As a consequence, we give a quick proof of the gelation phenomenon. Additionally, this connection leads to a series expression for the extinction probabilities of the multi-type Poisson branching process. While the correspondence highlighted here is fairly straightforward, it illuminates a deep connection between these two paradigms which we hope will continue to reveal new insights and potential for cross-disciplinary research.We begin by recalling some of the notation and concepts from the theory of vector multiplicative coalescent processes. For a detailed account of this theory, we refer the reader to [15, 16].
1.1. Vector multiplicative coalescent processes
Roughly speaking, a vector multiplicative coalescent process is a Markov process that describes cluster merger dynamics, where the weight of each cluster is a finite-dimensional vector. Following the convention in [16], to define this process formally, we use bra-ket notation. Specifically, $|{\bf x}\rangle$ will denote the column vector representation of vector ${\bf x}\in\mathbb{R}^{k}$, and $\langle{\bf x}|$ will denote the row vector representation of vector ${\bf x}\in\mathbb{R}^{k}$. For $c\in\mathbb{R}$ and ${\bf x}\in\mathbb{R}^{k}$, $\,c|{\bf x}\rangle$ will represent the product $c{\bf x}$, a column vector. Correspondingly, $\langle{\bf x}|{\bf y}\rangle=\langle{\bf y}|{\bf x}\rangle$ will be the dot product of ${\bf x}$ and ${\bf y}$ in $\mathbb{R}^{k}$. Finally, for a matrix $M\in\mathbb{R}^{k\times k}$, $\langle{\bf x}|M|{\bf y}\rangle$ will represent the product ${\bf x}^{\sf T}M{\bf y}$ resulting in a scalar.
@@ 1.1. Vector multiplicative coalescent processes
with initial conditions $\zeta_{\bf x}(0)=\alpha_{i}$ if $x={\bf e}_{i}$, and $\zeta_{\bf x}(0)=0$ otherwise. These equations describe the evolution of the size distribution of clusters in a system where clusters merge multiplicatively as defined in (1). More specifically, the first term on the right hand side in (2) corresponds to the depletion of clusters of size x after merging with other clusters and the second term corresponds to coagulation of smaller clusters to produce one of size x. For further details on Smoluchowski equations in various settings including applications, we refer the reader to [21, 18, 10, 19, 5, 14, 7].
1.2. The solution to the Smoluchowski equations
@@
with the same initial conditions coincide as long as the second order moments are convergent. The unique solution $\zeta_{\bf x}(t)$ of the above modified Smoluchowski equations was derived in [16]. Specifically, for a vector ${\bf x}\in\mathbb{Z}_{+}^{k}$ let $\,{\bf x}!=x_{1}!x_{2}!\ldots x_{k}!\,$ and for vectors ${\bf a}$ and ${\bf b}$ in $\mathbb{R}^{k}$ let $\,{\bf a}^{\bf b}=a_{1}^{b_{1}}a_{2}^{b_{2}}\ldots a_{k}^{b_{k}}\,$ whenever $a_{i}^{b_{i}}$ is uniquely defined for all $i$. Now, consider a complete graph $K_{k}$ consisting of vertices $\{1,\ldots,k\}$ with weights $w_{i,j}=w_{j,i}\geq 0$ assigned to its edges $[i,j]$ ($i\not=j$). Let the weight $W(\mathcal{T})$ of a spanning tree $\mathcal{T}$ be the product of the weights of all of its edges. Finally, let $\tau(K_{k},w_{i,j})=\sum\limits_{\mathcal{T}}W(\mathcal{T})$ denote the weighted spanning tree enumerator, i.e., the sum of weights of all spanning trees in $K_{k}$. In the case $w_{i,j}\equiv 1$, the weighted spanning tree enumerator simply equals the number of spanning trees of the graph. In terms of these notations, the solution to the modified Smoluchowski equations that yield the limiting fraction of clusters of size ${\bf x}$ is given by
@@
Outline of the paper.  The rest of the paper is organized as follows: Section 2 is devoted to a short proof of equation (4) based on the theory of random graphs, which could be of independent interest, as it provides a probabilistic perspective on the Smoluchowski equations. In Section 3, we discuss gelation and its connection to the multidimensional Lambert-Euler inversion. In Section 4, we recall some basic concepts from the theory of multi-type branching processes. We conclude by stating the connection between the vector-multiplicative coalescent process and multi-type Poisson branching processes and its applications.
2. Random graphs and an alternative proof of equation (4)
The following alternative proof uses the connection between the vector multiplicative coalescent process $C$ and the theory of random graphs. This connection works as follows: clusters of particles in $C$ correspond to small connected components of a certain random graph $G$. The so-called giant component of $G$ corresponds to the gel in $C$ for details, see the next section. The correspondence between $C$ and $G$ was first described in [16], to which we refer the reader for a full account. Here, we include only enough detail to describe the proof of (4).
@@ 2. Random graphs and an alternative proof of equation (4)
With $\alpha$ and $V$ fixed, there will be a critical value $t^{\prime}$ of $t$ below which there is almost surely no giant component, and above which there is almost surely a giant component. Denote by $D[\alpha_{i}]$ the diagonal matrix with the $\alpha_{i}$ on the diagonal. Then, this critical value $t^{\prime}$ is such that the largest eigenvalue of the matrix $t^{\prime}VD[\alpha_{i}]$ equals 1 (i.e., $t^{\prime}$ is the reciprocal of the spectral radius of $VD[\alpha_{i}]$).
@@
For simplicity, we first address the case where $k=\alpha_{1}=v_{11}=1$. In this case, setting ${\bf x}=l$, we have
(5) $\zeta_{l}(t)=\phi_{l}(t)=\frac{l^{l-2}t^{l-1}e^{-lt}}{l!}$
The random graphs proof of this is as follows. We calculate $p_{l}(t)=l\phi_{l}(t)$, which is the probability that a given vertex lies in a connected component $G_{l}$ of size $l$ (which, with probability tending to 1, is a tree - see [6] for a proof). Now, setting $p=t/n$ for the edge probability, we see that:
  • There are $\binom{n-1}{l-1}$ choices for the other vertices in $G_{l}$
  • There are $l^{l-2}$ labeled trees on these chosen vertices
  • Each tree is present with probability $p^{l-1}$
  • The edges between $G_{l}$ and the other vertices are missing with probability $(1-p)^{l(n-l)}$
@@
$p_{l}(t)$ $\sim\binom{n-1}{l-1}\cdot l^{l-2}\cdot p^{l-1}\cdot(1-p)^{l(n-l)}\sim\frac{n^{l-1}}{(l-1)!}\cdot l^{l-2}\cdot\left(\frac{t}{n}\right)^{l-1}\cdot e^{-pln}$
$=\frac{1}{(l-1)!}\cdot l^{l-2}\cdot t^{l-1}\cdot e^{-tl}=\frac{l^{l-1}t^{l-1}e^{-lt}}{l!},$
which, recalling that $p_{l}(t)=l\phi_{l}(t)$, establishes (5).
Next we consider the general case. For this we once again require the weighted spanning tree enumerator defined above, although here it arises in a slightly different form. Write $K_{\bf x}(V)$ for the complete graph on $\langle{\bf x}|\boldsymbol{1}\rangle$ vertices with $x_{i}$ vertices of type $i$, and where the weight of an edge between vertices of types $i$ and $j$ is $v_{ij}$. Suppressing the dependence on $V$, write $T_{\bf x}$ for the sum of the weights of all the spanning trees of $K_{\bf x}(V)$. Then, from Theorem 3.8 in [16],
$T_{\bf x}=\frac{\tau(K_{k},x_{i}x_{j}v_{i,j})}{{\bf x}^{\bf 1}}(V{\bf x})^{{\bf x}-{\bf 1}}.$
@@
  • These form a tree with probability $T_{\bf x}\cdot(t/n)^{\langle{\bf x}|\boldsymbol{1}\rangle-1}$
@@
$p_{\bf x}(t)$ $\sim\langle{\bf x}|\boldsymbol{1}\rangle\frac{n^{\langle{\bf x}|\boldsymbol{1}\rangle-1}\alpha_{1}^{x_{1}}\cdots\alpha_{k}^{x_{k}}}{x_{1}!\cdots x_{k}!}\cdot T_{\bf x}\cdot\left(\frac{t}{n}\right)^{\langle{\bf x}|\boldsymbol{1}\rangle-1}\cdot{\rm exp}\left(-\sum_{i}\sum_{j}x_{i}v_{ij}\alpha_{j}t\right)$
$=\langle{\bf x}|\boldsymbol{1}\rangle\frac{\alpha_{1}^{x_{1}}\cdots\alpha_{k}^{x_{k}}}{x_{1}!\cdots x_{k}!}\cdot T_{\bf x}\cdot t^{\langle{\bf x}|\boldsymbol{1}\rangle-1}\cdot{\rm exp}\left(-\sum_{i}\sum_{j}x_{i}v_{ij}\alpha_{j}t\right),$
@@
$\zeta_{\bf x}(t)=\phi_{\bf x}(t)=\frac{\alpha_{1}^{x_{1}}\cdots\alpha_{k}^{x_{k}}}{x_{1}!\cdots x_{k}!}\cdot T_{\bf x}\cdot t^{\langle{\bf x}|\boldsymbol{1}\rangle-1}\cdot{\rm exp}\left(-\sum_{i}\sum_{j}x_{i}v_{ij}\alpha_{j}t\right),$
@@
3. Gelation and Lamber-Euler Inversion
It is natural to ask if the total mass for a system described by the Smoluchowski coagulation model is conserved. In [13, 20, 3, 22, 23], it was shown that for certain classes of coagulation kernels, this is not always possible as a system can undergo a phase transition. For example, Hendriks et al. [11] studied this phase transition for the model with the multiplicative kernel (in one-dimensional setting), and they showed that the total mass decreases after a finite time, observing a surprising phenomenon known as gelation. Broadly speaking, gelation is the break down of conservation of mass in the system and can be physically interpreted as being caused by the formation of a giant cluster called the gel.
For the coalescent process defined in section 1.1, the initial total mass vector is assumed to be $\,\sum_{\bf x}\zeta_{\bf x}(0)|{\bf x}\rangle=|\boldsymbol{\alpha}\rangle$. Therefore, we define the gelation as the loss of total mass $\,\sum_{\bf x}\zeta_{\bf x}(t)|{\bf x}\rangle\,$ of the system after a critical time $T_{gel}$ called gelation time. That is
@@ 3. Gelation and Lamber-Euler Inversion
where $\rho(A)$ denotes the spectral radius of matrix $A$.
@@
The following theorem implies the existence of a unique solution to the multidimensional Lambert-Euler problem (see [16, Theorem 1.1]).
@@
Finally, in [16], it was shown that the total mass $\sum_{\bf x}\zeta_{\bf x}(t)|{\bf x}\rangle$ of the vector multiplicative coalescent process could be expressed in terms of the solution of the multidimensional Lambert-Euler inversion problem as follows:
@@
Then, Theorem 3.1 proves the existence of the gelation phenomenon for the vector multiplicative coalescent process and the gelation time given in (7). Specifically,
$\sum_{\bf x}\zeta_{\bf x}(t)|{\bf x}\rangle\left\{\begin{array}[]{ll}=|{\boldsymbol{\alpha}}\rangle&\mbox{if $t\leq 1/\rho\left(VD[\alpha_{i}]\right)$}\\ <|{\boldsymbol{\alpha}}\rangle&\mbox{if $t>1/\rho\left(VD[\alpha_{i}]\right)$}\end{array}\right.$
@@ 4. Multi-Type Poisson Branching Processes
In other words, the entry $m_{ij}$ represents the average number of offspring of type $j$ produced by an individual of type $i$ in one generation. In the case where the entries $m_{i,j}$ vary, we have a multi-type inhomogenous branching process.
@@ 4.1. Extinction Probability
As gelation is to vector multiplicative coalescent processes, a fundamental question in the study of branching processes is whether the process becomes extinct as the number of generations $n$ tends to infinity. Recall that in the classical case, if the expected number of the offspring distribution is less than (or equal to) 1, then eventually the population dies out. But if the expected number is greater than 1, then the probability of extinction is the smallest non-negative solution to the equation $s=\varphi(s)$, where $\varphi$ is the probability generating function of the corresponding offspring distribution.